Let A and B be two players who are playing the game to hit the target. The probabilities of hitting the target by A and B is and , respectively. What is the probability that exactly one of them hit the target?
Show answer & explanation
Correct answer: C
Explanation
The correct answer is option 3 i.e. .
P(A) = 2/3
P(B) = 3/4
P(A') = 1 - P(A) = 1 - (2/3) = 1/3
P(B') = 1 - P(B) = 1 - (3/4) = 1/4
There are two scenarios:
A hits and B misses (A and B') and A misses and B hits (A' and B)
P(A and B') = P(A) × P(B') = (2/3) × (1/4) = 2/12 = 1/6
P(A' and B) = P(A') × P(B) = (1/3) × (3/4) = 3/12 = 1/4
P(exactly one hits) = P(A and B') + P(A' and B) = (1/6) + (1/4) = 5/12
Therefore, the probability that exactly one of them hits the target is 5/12.